# Effective Bertini theorems and zeros of $p$-adic forms of degree 7

### v1

@@ Abstract.
We establish an effective Bertini-type theorem for hypersurfaces $X_{f}\colon f=0$ defined over a finite field $k$ for which $f$ has no linear factors over the algebraic closure $\overline{k}$. Given a line $L$ defined over $k$ and a nonreduced $\overline{k}$-point $x$ on $X_{f}\cap L$, we give an upper bound on the number of planes $P$ containing $L$ for which $X_{f}\cap P$ contains a line through $x$. Underlying this result is a factorization algorithm for bivariate polynomials originally due to Kaltofen, which we present with slightly relaxed hypotheses. Our primary application is to Artin’s conjecture on $p$-adic forms of degree 7: if $K/\mathbb{Q}_{p}$ is a finite extension with residue field isomorphic to $\mathbb{F}_{q}$ and $F(x_{0},\ldots,x_{n})\in K[x_{0},\ldots,x_{49}]$ is a homogeneous form of degree 7, then there exists a $K$-solution to $F=0$ whenever q>679q>{679}. This improves on a result of Wooley.
@@ 1. Introduction
In this work, we prove an effective Bertini-type theorem with a focus on ruling out certain factorizations in planar slices of a hypersurface in $\mathbb{P}^{n}$. Over a finite field $\mathbb{F}_{q}$ for $q$ sufficiently large, this ensures existence of a plane with the desired properties. We give an application of this to a conjecture of Artin concerning zeros of degree seven forms over $p$-adic fields.
@@ 1.1. On Bertini theorems
In another direction, Kaltofen [Kal95] gives an effective Bertini theorem replacing smoothness with geometric irreducibility for planar slices: given a geometrically irreducible hypersurface $X\subset\mathbb{P}^{n}$ for $n\geq 3$, he gives an upper bound on number of planes $P\subset\mathbb{P}^{n}$ for which $X\cap P$ is not geometrically irreducible in terms of the degree of $X$. Over a finite field $\mathbb{F}_{q}$, this is essentially an effective version of the Hilbert Irreducibility Theorem. Cafure and Matera [CM06] modified this approach further to bound how often such a planar slice contains irreducible components of small degree. We give a modest improvement on this result below.
Theorem A (See Theorem 2.1 and [CM06, Corollaries 3.2, 3.4]).
Suppose $f(x_{1},\ldots,x_{n})\in\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ is an absolutely irreducible (not necessarily homogeneous) polynomial of total degree $d$. Define a family of bivariate polynomials
$f_{\mathbf{v},\mathbf{w},\mathbf{z}}=f(X+v_{1},w_{2}X+z_{2}Y+v_{2},\ldots,w_{n}X+z_{n}Y+v_{n})\in\mathbb{F}_{q}[X,Y]$
for
$(\mathbf{v},\mathbf{w},\mathbf{z})=(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n},z_{2},\ldots,z_{n})\in\mathbb{F}_{q}^{3n-2}.$
@@
    There exist at most
    $\frac{1}{8}\left(3d^{4}-2d^{3}+13d^{2}+2d\right)q^{3n-3}$
    tuples $(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}$ such that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ is not absolutely irreducible.
@@
    Fix a positive integer $D<d$. Then there exist at most
    $\frac{d}{8}\left(-D^{4}+4D^{3}d-6D^{3}+12D^{2}d-11D^{2}+8Dd-6D+16d\right)q^{3n-3}$
    tuples $(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}$ for which $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ has an irreducible factor of degree at most $D$ over $\overline{\mathbb{F}_{q}}$.
We modify Theorem A further to address the situation where $f$ may be reducible over $\overline{\mathbb{F}_{q}}$, but possesses no linear factors, subject to the hypothesis that for some $(\mathbf{v}_{0},\mathbf{w}_{0})\in\mathbb{F}_{q}^{2n-1}$ we have $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}(X,0)$ has a simple root at $\alpha_{0}\in\overline{\mathbb{F}_{q}}$. More precisely, we construct a polynomial $\Psi(Z_{2},\ldots,Z_{n})$ which, if nonzero, implies that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}$ has no linear factor vanishing at $(\alpha_{0},0)$; for the full statement, see Theorem 2.6. In particular, we obtain the following corollary.
Theorem B (See Corollary 2.7).
Suppose $f(x_{1},\ldots,x_{n})$ is a (not necessarily homogeneous) polynomial of total degree $d$ over $\mathbb{F}_{q}$, possessing no linear factors over $\overline{\mathbb{F}_{q}}[x_{1},\ldots,x_{n}]$. Moreover, suppose there exist $(\mathbf{v}_{0},\mathbf{w}_{0})\in\mathbb{F}_{q}^{2n-1}$ such that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}(X,0)\in\overline{\mathbb{F}_{q}}[X]$ is degree $d$ with a simple root $\alpha_{0}\in\overline{\mathbb{F}_{q}}$.
Then there exist at most $(3d-3)q^{n-2}$ tuples $\mathbf{z}\in\mathbb{F}_{q}^{n-1}$ such that $f_{\mathbf{v}_{0},\mathbf{w}_{0},\mathbf{z}}$ has no linear factor over $\overline{\mathbb{F}_{q}}$ containing $(\alpha_{0},0)$.
Our main application is to Artin’s conjecture for $p$-adic forms of degree 7.
@@ 1.2. On Artin’s conjecture
Let $d>3$ be prime and $K/\mathbb{Q}_{p}$ be a finite extension. Then for all $n\geq d^{2}$, every degree $d$ form $F\in K[x_{0},\ldots x_{n}]$ has a zero in $K$, i.e. $K$ is $C_{2}(d)$.
In degree $d=7$, we improve the best known effective lower bound on $q$ for which Artin’s conjecture holds.
Theorem C.
Let $K/\mathbb{Q}_{p}$ be a $p$-adic field with residue field $\mathbb{F}_{q}$. If q>679q>{679}, then for all $n\geq 49$, every degree 7 form $F\in K[x_{0},\ldots,x_{n}]$ has a zero in $K$, i.e. $K$ is $C_{2}(7)$.
To establish Theorem C, we follow a $p$-adic minimization approach due to Birch and Lewis [BL59] and later refined by Laxton and Lewis [LL65] to give effective bounds on $q$ for which Artin’s conjecture holds in degrees $d=5,7,11$. They proposed to reduce to the case of so-called reduced forms $F$, then use an effective version of the Lang–Weil bound for the number of $\mathbb{F}_{q}$-points on a variety to produce a nonsingular zero of $\overline{F}$, the image of $F$ under the residue map, when $q$ is sufficiently large. This then lifts to a $K$-zero of $F$ by Hensel’s lemma.
Refinements of this general strategy have shown that $K$ is $C_{2}(5)$ whenever q>5q>{5} due to [LY96, HB10, Dum17] and to forthcoming work of the authors, and that $K$ is $C_{2}(7)$ (resp. $C_{2}(11)$) whenever q>883q>{883} (resp. q>8053q>{8053}) due to [Kna01, Woo08]. This strategy breaks down for prime degrees $d>11$, since $d$ can be written as the sum of composite numbers, which allows for the hypersurface given by $\overline{F}=0$ to have only totally nonreduced nonlinear components.
@@
Inspired by the work of Leep and Yeomans in the quintic case, we further exploit the structure of reduced degree 7 forms either to find nonsingular zeros of $\overline{F}$, or to show that (a dehomogenization of) $\overline{F}$ satisfies the hypotheses of Theorem B. Upon applying Theorem B, we produce a plane curve with at least two singular $\mathbb{F}_{q}$-points and a geometrically irreducible component; these extra singularities lower the $q$ threshold for which we can find nonsingular points on the curve, and hence zeros of $F$ in $K$.
@@ 1.3. Notation and organization
This paper is organized as follows. In Section 2 we describe a factorization algorithm for bivariate forms, which we use to prove Theorem B. We give a proof of Theorem A in Section 2.3. In Section 3 we record some facts about degree 7 forms over $\mathbb{F}_{q}$ and in Section 4 we recall the definition of reduced forms and collect auxiliary results about them. Finally, we prove Theorem C in Section 5.
@@ 2. Effective Bertini theorems
Let $k$ be an arbitrary field and suppose $f(x_{0},\ldots,x_{n})$ is a homogeneous polynomial of degree $d$ over $k$, irreducible over $\overline{k}$. Recall that generically, the intersection of the projective hypersurface $X_{f}\colon f=0$ remains geometrically irreducible when intersected with a hyperplane; this is Bertini’s theorem for irreducibility. However, when $k=\mathbb{F}_{q}$, this does not suffice to guarantee the existence of even one such hyperplane.
By an effective Bertini theorem, we refer to a result which bounds how often the non-generic behavior occurs. When $k=\mathbb{F}_{q}$, this can then be used give a lower bound on $q$ for which the generic behavior is guaranteed to occur at least once. We will be particularly interested in intersecting $X_{f}$ with a plane, rather than a hyperplane, and find it convenient to work in affine coordinates.
@@
$f_{\mathbf{v},\mathbf{w},\mathbf{z}}=f(1,X+v_{1},w_{2}X+z_{2}Y+v_{2},\ldots,v_{1},w_{n}X+z_{n}Y+v_{n})\in k[X,Y]$
@@
$(\mathbf{v},\mathbf{w},\mathbf{z})=(v_{1},\ldots,v_{n},w_{2},\ldots,w_{n},z_{2},\ldots,z_{n})\in k^{3n-2}.$
We think of $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ as cutting out an affine plane curve, obtained by slicing the $x_{0}\neq 0$ affine patch of $X_{f}$ by the plane determined by $(\mathbf{v},\mathbf{w},\mathbf{z})$. If the original $f(x_{0},\ldots,x_{n})$ is absolutely irreducible, then generically so is $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$.
This was made effective by Kaltofen, using an algorithm for testing irreducibility of a bivariate polynomial to deduce the existence of a polynomial condition on $(\mathbf{v},\mathbf{w},\mathbf{z})$ whose vanishing is implied by reducibility of $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ [Kal95]. Cafure and Matera modified this approach to detect how often $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ has factors of small degree [CM06]. We recall Theorem A, a mild strengthening of [CM06, Corollaries 3.2, 3.4]. For completeness, we describe the necessary modifications to the proofs of [CM06, Corollaries 3.2, 3.4] in §2.3.
Theorem 2.1 (See [CM06, Corollaries 3.2, 3.4]).
Suppose $f(x_{1},\ldots,x_{n})\in\mathbb{F}_{q}[x_{1},\ldots,x_{n}]$ is an absolutely irreducible (not necessarily homogeneous) polynomial of total degree $d$.
@@
    There exist at most
    $\frac{1}{8}\left(3d^{4}-2d^{3}+13d^{2}+2d\right)$
    tuples $(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}$ such that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ is not absolutely irreducible.
@@
    Fix a positive integer $D<d$. Then there exist at most
    $\frac{d}{8}\left(-D^{4}+4D^{3}d-6D^{3}+12D^{2}d-11D^{2}+8Dd-6D+16d\right)q^{3n-3}$
    tuples $(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}$ for which $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ has an irreducible factor of degree at most $D$ over $\overline{\mathbb{F}_{q}}$.
@@
Let $d=7$ and assume $f(x_{1},\ldots,x_{n})$ is irreducible over $\overline{\mathbb{F}_{q}}$. By Theorem 2.1(i), when q>896q>{896} we have that there must exist at least one $(\mathbf{v},\mathbf{w},\mathbf{z})\in\mathbb{F}_{q}^{3n-2}$ such that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}(X,Y)$ is absolutely irreducible. Suppose instead that we only require that $f_{\mathbf{v},\mathbf{w},\mathbf{z}}(X,Y)$ does not have a linear factor over $\overline{\mathbb{F}_{q}}$. Then applying Theorem 2.1(ii) with $D=1$, such a $(\mathbf{v},\mathbf{w},\mathbf{z})$ exists whenever q>224q>{224}.
The latter condition corresponds to the plane curve cut out by $f_{\mathbf{v},\mathbf{w},\mathbf{z}}$ having no linear irreducible components over $\overline{\mathbb{F}_{q}}$. Since $d=7$, this forces at least one of the irreducible components to also have multiplicity 1 and be defined over $\mathbb{F}_{q}$. This was a key insight used by Wooley [Woo08] to show that $p$-adic fields are $C_{2}(7)$ when q>883q>{883}. (resp. $C_{2}(11)$ when q>8053q>{8053}).
We will need a variant of Theorem 2.1 describing how often $X_{f}$ has linear factors when sliced by planes through a specified line $L$, under some hypotheses on $X_{f}\cap L$. Before stating and proving this variant in Corollary 2.7, we briefly describe a relaxation of Kaltofen’s factorization algorithm for $f(X,Y)$ over $\overline{k}$, which gives partial information when $f(X,0)$ is not squarefree.

### v2

# Effective Bertini theorems and zeros of $p$-adic forms of degrees 7 and 11

@@ Abstract.
We establish an effective Bertini-type theorem for hypersurfaces $X_{f}\colon f=0$ defined over a finite field $k$ for which $f$ has no linear factors over the algebraic closure $\overline{k}$. Given a line $L$ defined over $k$ and a nonreduced $\overline{k}$-point $x$ on $X_{f}\cap L$, we give an upper bound on the number of planes $P$ containing $L$ for which $X_{f}\cap P$ contains a line through $x$. Underlying this result is a factorization algorithm for bivariate polynomials originally due to Kaltofen, which we present with slightly relaxed hypotheses. Our primary application is to Artin’s conjecture on $p$-adic forms of prime degree $d$: if $K/\mathbb{Q}_{p}$ is a finite extension with residue field isomorphic to $\mathbb{F}_{q}$ and $F\in K[x_{0},\ldots,x_{d^{2}}]$ is homogeneous of degree $d$, the conjecture states $F$ has a nontrivial zero in $K$. We show this conjecture holds whenever q>679q>{679} for $d=7$ and q>7393q>{7393} for $d=11$, improving upon a result of Wooley.
@@ 1. Introduction
In this work, we prove an effective Bertini-type theorem with a focus on ruling out certain factorizations in planar slices of a hypersurface in $\mathbb{P}^{n}$. Over a finite field $\mathbb{F}_{q}$ for $q$ sufficiently large, this ensures existence of a plane with the desired properties. We give an application of this to a conjecture of Artin concerning zeros of forms over $p$-adic fields.
@@ 1.1. On Bertini theorems
In another direction, Kaltofen [KAL95] gave an effective Bertini theorem replacing smoothness with geometric irreducibility for planar slices: given a geometrically irreducible hypersurface $X\subset\mathbb{P}^{n}$ for $n\geq 3$, he gave an upper bound on number of planes $P\subset\mathbb{P}^{n}$ for which $X\cap P$ is not geometrically irreducible in terms of the degree of $X$. Over a finite field $\mathbb{F}_{q}$, this is essentially an effective version of the Hilbert Irreducibility Theorem. Cafure and Matera [CM06] modified this approach further to bound how often such a planar slice contains irreducible components of small degree. We give a modest improvement on these results and extend to reducible hypersurfaces (see Theorem 2.1). We also give a sharper result if one is interested only in the existence of a single such plane.
Theorem A (effective Bertini irreducibility for planes, existence form (see Theorem 2.2)).
Let $d\geq 2$, $n\geq 3$, and $q>{2d^{2}}$. Suppose $f\in\mathbb{F}_{q}[x_{0},\ldots,x_{n}]$ is a degree $d$ form, irreducible over $\overline{\mathbb{F}_{q}}$, and $X_{f}\colon f=0$ is the associated geometrically integral hypersurface in $\mathbb{P}^{n}$ defined over $\mathbb{F}_{q}$.
@@
    If $q>{{\frac{d}{8}\left(3d^{3}-2d^{2}-3d+2\right)}}$, there exists a plane $P\subset\mathbb{P}^{n}$ defined over $\mathbb{F}_{q}$ such that $X_{f}\cap P$ is a geometrically integral plane curve of degree $d$, defined over $\mathbb{F}_{q}$.
@@
    Fix an integer $0<D<d$. If
    $q>\frac{dD}{8}\left(-D^{3}+4dD^{2}-6D^{2}+12dD-11D+8d-6\right),$
    there exists a plane $P\subset\mathbb{P}^{n}$ defined over $\mathbb{F}_{q}$ such that $X_{f}\cap P$ is a plane curve of degree $d$, defined over $\mathbb{F}_{q}$, with no irreducible components of degree at most $D$ defined over $\overline{\mathbb{F}_{q}}$.
A key technical tool is a more flexible version of [KAL95, Theorem 5] and [CM06, Theorem 3.3], which we briefly describe geometrically. Given a geometrically integral hypersurface $X_{f}$ of degree $d$ and a line $L$ meeting $X_{f}$ transversely in at least one marked geometric point, among planes $P$ containing $L$, $X_{f}\cap P$ is generically a geometrically integral plane curve of degree $d$. In Theorem 2.9, we parametrize the planes $P$ and describe polynomial conditions which must be satisfied if $X_{f}\cap P$ has components of small degree containing the marked point. We use Theorem 2.9 not only as an intermediate step in the proof of Theorem A, but also find an application to Artin’s conjecture for $p$-adic forms of degrees 7 and 11.
@@ 1.2. On Artin’s conjecture
Let $K/\mathbb{Q}_{p}$ be a finite extension and $d>3$ prime. For all $n\geq d^{2}$, every degree $d$ form $F\in K[x_{0},\ldots x_{n}]$ has a nontrivial zero in $K$, i.e. $K$ is $C_{2}(d)$.
In degrees $d=7$ and $d=11$, we improve the best known effective lower bound on $q$ for which Artin’s conjecture holds.
Theorem B.
Let $K/\mathbb{Q}_{p}$ be a $p$-adic field with residue field $\mathbb{F}_{q}$.
  1. (i)
    If q>679q>{679} then every degree $7$ form $F\in K[x_{0},\ldots x_{n}]$ has a nontrivial zero in $K$, i.e. $K$ is $C_{2}(7)$.
  2. (ii)
    If q>7393q>{7393} then every degree $11$ form $F\in K[x_{0},\ldots x_{n}]$ has a nontrivial zero in $K$, i.e. $K$ is $C_{2}(11)$.
To establish Theorem B, we follow a $p$-adic minimization approach due to Birch and Lewis [BL59] and later refined by Laxton and Lewis [LL65] to give effective bounds on $q$ for which Artin’s conjecture holds in degrees $d=5,7,11$. They proposed to reduce to the case of so-called reduced forms $F$, then use an effective version of the Lang–Weil bound for the number of $\mathbb{F}_{q}$-points on a variety to produce a nonsingular zero of $\overline{F}$, the image of $F$ under the residue map, when $q$ is sufficiently large. This then lifts to a $K$-zero of $F$ by Hensel’s lemma. We recall the definition of reduced forms and basic facts about them in §4.1.
Refinements of this general strategy have shown that $K$ is $C_{2}(5)$ whenever q>5q>{5} due to [LY96, HEA10, DUM17] and to forthcoming work of the authors, and that $K$ is $C_{2}(7)$ (resp. $C_{2}(11)$) whenever q>883q>{883} (resp. q>8053q>{8053}) due to [KNA01, WOO08]. This strategy breaks down for prime degrees $d>11$, since such $d$ can be written as the sum of composite numbers, which allows for the hypersurface given by $\overline{F}=0$ to have only totally nonreduced nonlinear components.
@@
Inspired by the work of Leep and Yeomans in the quintic case, we further exploit the structure of reduced forms of degrees 7 and 11 to either find nonsingular zeros of $\overline{F}$, or to show that $\overline{F}$ has two (necessarily singular) $\mathbb{F}_{q}$-zeros spanning a line $L$ meeting $\overline{X_{F}}$ transversely at a point. The latter is amenable to an application of our modified effective Bertini result, Theorem 2.9, producing a plane curve with at least two singular $\mathbb{F}_{q}$-points and a geometrically irreducible component; these extra singularities lower the $q$ threshold for which we can find nonsingular points on the curve, and hence zeros of $F$ in $K$.
@@ 1.3. Notation and organization
This paper is organized as follows. In Section 2 we discuss Bertini theorems for irreducibility, stating two versions in Theorems 2.1 and 2.2, the latter of which specializes to Theorem A. In §2.1 we describe a factorization algorithm for bivarate polynomials, which we employ in §2.2 to prove Theorem 2.9. In Section 3 we record some facts about degree 7 and 11 forms over $\mathbb{F}_{q}$, while in Section 4 we recall the definition of reduced forms and prove some consequences in the degree 7 and 11 cases. Finally, we prove Theorem B in Section 5.
@@ 2. Effective Bertini theorems
Let $k$ be an arbitrary field, $n\geq 3$, and suppose $f\in k[x_{0},\ldots,x_{n}]$ is a degree $d$ form defined over $k$, irreducible over $\overline{k}$. Recall that generically, the intersection of the projective hypersurface $X_{f}\colon f=0$ with a hyperplane remains geometrically irreducible; this is Bertini’s theorem for irreducibility. However, when $k=\mathbb{F}_{q}$, this does not suffice to guarantee the existence of even one such hyperplane.
By an effective Bertini theorem, we refer to a result which bounds how often the non-generic behavior occurs. When $k=\mathbb{F}_{q}$, the approach can be adapted to give a lower bound on $q$ for which the generic behavior is guaranteed to occur at least once. We will be particularly interested in intersecting $X_{f}$ with a plane, rather than a hyperplane, and find it convenient to work in affine coordinates.
@@
$f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}=f(1,X+\beta_{1},\gamma_{2}X+\delta_{2}Y+\beta_{2},\ldots,\beta_{1},\gamma_{n}X+\delta_{n}Y+\beta_{n})\in k[X,Y]$
@@
$(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})=(\beta_{1},\ldots,\beta_{n},\gamma_{2},\ldots,\gamma_{n},\delta_{2},\ldots,\delta_{n})\in k^{3n-2}.$
One may view $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ as cutting out an affine plane curve, obtained by slicing the $x_{0}\neq 0$ affine patch of $X_{f}$ by the plane determined by $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$. If the original $f(x_{0},\ldots,x_{n})$ is absolutely irreducible, then generically so is $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$.
More generally, if $f$ factors as $f=\prod_{1\leq i\leq t}f_{i}^{e_{i}}$ for $f_{i}$ distinct and irreducible and $e_{i}\geq 1$, we say $f$ has factorization type $(d_{1}^{e_{i}},\ldots,d_{t}^{e_{t}})$. If $x_{0}\nmid f$, then generically $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ shares the same factorization type as $f$. For brevity, if $e_{i}=1$, we omit it from the type notation. Note that the factorization type is not invariant under base extension; for the most part we will be interested in the factorization type over $\overline{k}$.
For $f$ irreducible over $\overline{k}$, Kaltofen used an algorithm for testing irreducibility of a bivariate polynomial to deduce polynomial conditions on $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$ whose vanishing is implied by reducibility of $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ [KAL95]. Cafure and Matera modified this approach to detect how often $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has factors of small degree [CM06]. We offer an extension of their results in Theorem 2.1 below, which specializes to a mild strengthening of [CM06, Corollaries 3.2, 3.4] in the $t=1$ case.
Theorem 2.1 (effective Bertini irreducibility for planes).
Let $d\geq 2$ and $n\geq 3$. Suppose $f\in\mathbb{F}_{q}[x_{0},\ldots,x_{n}]$ is a degree $d$ form with factorization type $(d_{1}^{e_{1}},\ldots,d_{t}^{e_{t}})$ over $\overline{\mathbb{F}_{q}}$ such that $x_{0}\nmid f$.
@@
    For all but
    $\left(2\left(\sum_{1\leq i\leq t}d_{i}\right)^{2}+\sum_{1\leq i\leq t}\frac{d_{i}}{8}\left(3d_{i}^{3}-2d_{i}^{2}-3d_{i}+2\right)\right)q^{3n-3}$
    tuples $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})\in\mathbb{F}_{q}^{3n-2}$, $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,Y)$ has factorization type $(d_{1}^{e_{1}},\ldots,d_{t}^{e_{t}})$ over $\overline{\mathbb{F}_{q}}$.
@@
    Fix $0\leq D_{i}<d_{i}$ for $1\leq i\leq t$. For all but
    $\left(2\left(\sum_{1\leq i\leq t}d_{i}\right)^{2}+\sum_{1\leq i\leq t}\frac{d_{i}D_{i}}{8}\left(-D_{i}^{3}+4d_{i}D_{i}^{2}-6D_{i}^{2}+12d_{i}D_{i}-11D_{i}+8d_{i}-6\right)\right)q^{3n-3}$
    tuples $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})\in\mathbb{F}_{q}^{3n-2}$, $(f_{i})_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,Y)$ has no nonconstant factor of degree at most $D_{i}$ over $\overline{\mathbb{F}_{q}}$ for all $i$.
A consequence of Theorem 2.1 is an effective bound on $q$ for which the existence of at least one $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$ for which $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has the same factorization type as $f$ is guaranteed (and similarly for $f$ lacking factors of small degree). We prove a stronger existence result directly, under an additional hypothesis on $q$.
Theorem 2.2 (effective Bertini irreducibility for planes, existence form).
Let $d\geq 2$, $n\geq 3$, and $q>{2d^{2}}$. Suppose $f\in\mathbb{F}_{q}[x_{0},\ldots,x_{n}]$ is a degree $d$ form with factorization type $(d_{1}^{e_{1}},\ldots,d_{t}^{e_{t}})$ over $\overline{\mathbb{F}_{q}}$ such that $x_{0}\nmid f$.
  1. (i)
    If
    $q>\sum_{1\leq i\leq t}\frac{d_{i}}{8}\left(3d_{i}^{3}-2d_{i}^{2}-3d_{i}+2\right),$
    there exists $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})\in\mathbb{F}_{q}^{3n-2}$ such that $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has factorization type $(d_{1}^{e_{1}},\ldots,d_{t}^{e_{t}})$ over $\overline{\mathbb{F}_{q}}$.
  2. (ii)
    Fix $0\leq D_{i}<d_{i}$ for $1\leq i\leq t$. If
    $q>\sum_{1\leq i\leq t}\frac{d_{i}D_{i}}{8}\left(-D_{i}^{3}+4d_{i}D_{i}^{2}-6D_{i}^{2}+12d_{i}D_{i}-11D_{i}+8d_{i}-6\right),$
    there exists $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})\in\mathbb{F}_{q}^{3n-2}$ such that $(f_{i})_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has no nonconstant factor of degree at most $D_{i}$ over $\overline{\mathbb{F}_{q}}$ for all $i$.
Before proving Theorems 2.1 and 2.2, we establish some intermediates. In §2.1 we adapt Kaltofen’s bivariate factorization algorithm, relaxing what amounts to a requirement that $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,0)$ is squarefree. In §2.2, we use this to give a polynomial on $\boldsymbol{\delta}$ which vanishes when $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,Y)$ has small factors through a fixed (geometric) point. The proofs of Theorems 2.1 and 2.2 are then given in §2.3.
We conclude with some examples illustrating Theorems 2.1 and 2.2 in the degree 7 case.
@@
Let $d=7$, $n\geq 3$, and assume $f(x_{0},\ldots,x_{n})$ is irreducible over $\overline{\mathbb{F}_{q}}$. By Theorem 2.1(i), $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,Y)$ is irreducible over $\overline{\mathbb{F}_{q}}$ for all but $896q^{3n-3}$ choices of $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})\in\mathbb{F}_{q}^{3n-2}$.
If we merely need existence of at least one such $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$, we see that q>896q>{896} suffices. However, we can do slightly better using Theorem 2.2(i), which states if q>798q>{798}, there exists at least one $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})\in\mathbb{F}_{q}^{3n-2}$ such that $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ is irreducible over $\overline{\mathbb{F}_{q}}$.
Example 2.4.
Let $d=7$, $n\geq 3$, and assume $f(x_{0},\ldots,x_{n})$ is irreducible over $\overline{\mathbb{F}_{q}}$, as in Example 2.3. Suppose instead that we only require that $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}(X,Y)$ does not have a linear factor over $\overline{\mathbb{F}_{q}}$. Then applying Theorem 2.1(ii) with $D=1$, for all but $224q^{3n-3}$ choices of $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})\in\mathbb{F}_{q}^{3n-2}$, we have $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ has no linear factors. Using Theorem 2.2(ii), if q>126q>{126} then there exists at least one such $(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})$.
Example 2.5.
Let $d=7$, $n\geq 3$, and suppose $f(x_{0},\ldots,x_{n})$ has factorization type $(d_{1}^{e_{1}},\ldots,d_{t}^{e_{t}})$ over $\overline{\mathbb{F}_{q}}$. Let $N_{(d_{i}^{e_{i}})}$ count the number of $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ with the same type as $f$,
$N_{(d_{i}^{e_{i}})}=\#\left\{(\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta})\in\mathbb{F}_{q}^{3n-2}:f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}\text{ has type }(d_{1}^{e_{1}},\ldots,d_{t}^{e_{t}})\right\}.$
We can use Theorem 2.1(i) to give an upper bound for $\mathbb{F}_{q}^{3n-2}-N_{(d_{i}^{e_{i}})}$, and Theorem 2.2(i) to give a lower bound on $q$ for which at least one $f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ exists with type $(d_{1}^{e_{1}},\ldots,d_{t}^{e_{t}})$. This is illustrated in Table 2.1 for types not containing a linear factor defined over $\overline{\mathbb{F}_{q}}$.
Table 2.1. Illustrating Theorems 2.1(i) and 2.2(i) in degree 7
Factorization type of $f$ $(d_{1}^{e_{1}},\ldots,d_{t}^{e_{t}})$ $\mathbb{F}_{q}^{3n-2}-N_{(d_{i}^{e_{i}})}$ (Theorem 2.1(i)) $\exists f_{\boldsymbol{\beta},\boldsymbol{\gamma},\boldsymbol{\delta}}$ with type $(d_{i}^{e_{i}})$ (Theorem 2.2(i))
(7) $\leq 896q^{3n-3}$ q>798q>{798}
(5,2) $\leq 296q^{3n-3}$ q>198q>{198}
(4,3) $\leq 194q^{3n-3}$ q>98q>{98}
(3,2,2) $\leq 125q^{3n-3}$ q>98q>{98}
[…]